Two firms produce differentiated products with demand curves

p1 = a – q1 – bq2 and p2 = a – q2 – bq1. They both face constant average and marginal cost c and their profit functions are profit = (p1 – c)q1 and profit = (p2 — c)q2, respectively.

Solve the Bertrand game.

Hint: You need to solve the system of equations p1 = a − q1 − bq2 and p2 = a − q2 − bq1 for q1 and q2, which will allow you to write each firm’s profit as a function of prices and its marginal cost.
Answer #1

Page 1 (using ean:0 ; from 2 P₂ = a-qa-bq, %= a-bq, - p2 - 0 from 1 : Pi= a -9,- bqz > P, = 2 -9,- 6 (a-b9,-P2) * P, = Q -9,-

b=e- (et + b (tege) - face atob-ab - 1tb b(bp 1-62 Pinz atabat - b(02=P») – P2 235 - Corel) + ] 935 - ( 54 best 6-bit 2 bp₂-b

Solving W&@ simultaneously, Page 3 La featu]+ ct bo facebọcren)

P = Pa (1-6) + Qc + bal1-6) + bet be Page 4 too - → up, -b3p, = all-b) [a+b] + c(2+b] P. (22–62) - +b[ a[1-6] +c] = P, = (+6)

**If you liked the answer, then please upvote. Would be motivating for me. Thanks